OpenStudy (anonymous):

What is the solution to the system of equations? 2x – y = 7 y = 2x + 3 (2, 3) (2, 7) no solution infinite number of solutions

OpenStudy (anonymous):

-you can draw the two lines and see if they intersected or not, any intersection point is a solution for the system if there is no intersection points (the two lines are parallel) then there is no solution for this system

OpenStudy (anonymous):

no solution. 2 x - (3 + 2 x) = 7 -3 != 7